please help me slove this question .

the question : Measurement: use the form C= 5 on 9(f-32) to transform Temperatures from celicios (C) to fahrenheit (F) . write the transformation form from fahrenheit to celicios ?

Answer by Theo(13342)

formulas are:

F = (9/5)*C + 32

C = (F-32)*(5/9)

If C = 0, then F = 32

If C = 100, then F = 212

Going from C to F is ok.

If F = 32, C = 0

If F = 212, C = 100

Going from F to C is ok.

If C = 5, then F = (9/5)*C + 32 = 9 + 32 = 41 degrees farenheit.

If F = 41, then C = (F-32)*(5/9) = (41-32)*(5/9) = 9*(5/9) = 5 degrees Celsius.